# Efficient Generation of Grids and Traversal Graphs in Compositional Spaces towards Exploration and Path Planning Exemplified in Materials

### v1

@@
Efficient Generation of Grids and Traversal Graphs in Compositional Spaces towards Exploration and Path Planning Exemplified in Materials
@@ 1.1 Compositional Spaces
This work introduces a new high-performance library called nimplex or NIM library for simPLEX spaces, created exclusively for working with such spaces. It was written in low-level Nim language, allowing for careful optimizations, and then compiled with a native Python interface for general use. It provides an efficient implementation of (a) existing methods from literature (see Sec. 2.1 and 3.1), (b) modifications of existing methods (see Sec. 3.2), and (c) entirely new capabilities developed in this paper (see Sec. 4).
Neither compositional space nor nimplex is exclusive to any discipline; however, to better showcase its capabilities, two complex, highly-dimensional materials-related problems of high impact are highlighted. At the same time, translating them and their solutions to other areas, such as economics, can be done directly and is discussed.
@@ 1.2 Compositionally Complex Materials
An exemplar of how tackling highly-dimensional problems allows researchers to unlock novel solutions is the class of Compositionally Complex Materials (CCMs), which includes several sub-classes, such as Multi Principle Element Alloys (MPEAs), High Entropy Alloys (HEAs), High Entropy Ceramics (HECs), and High Entropy Metallic Glasses (HEMGs). CCMs are materials with at least several elements in significant fractions and was initiated by two pioneering 2004 works on HEAs by Yeh et al. [1] and by Cantor et al. [2], who independently proposed that equimolar (equal fractions) alloys with more than 5 (Yeh) or between 6 and 9 (Cantor) elements, could form single solid solutions (SSS) thanks to the configurational entropy stabilizing them. Other notable definitions include all materials with idealized configurational entropy $\Delta S_{conf}\geq R\ln{5}=1.61R$ [3] ($\approx 2.32$ bits of information in the composition x) or $\Delta S_{conf}\geq 1R$ [4] ($\approx 1.44$ bits).
@@
Regardless of the exact definition, while individual CCMs contain a few components, they always occupy very high dimensional problem spaces relative to other materials because they are not as restricted in terms of which elements are present. This results in homogeneous datasets occupying over 30-dimensional spaces (or 10-20 for specific problems, like refractory HEA [4]), which are orders of magnitude larger compared to traditional alloys with one or two primary elements. This introduces opportunities for finding exceptional alloys in little-explored chemical spaces, as demonstrated by some cases of excellent hardness [5], ductility [6], room temperature strength [7], and refractory strength [8, 9].
@@
In recent years, high-throughput thermodynamics-driven combinatorial studies on CCMs have been successfully performed to generate high-performance materials [10, 11], utilizing CALPHAD thermodynamic databases for CCMs/HEAs (e.g., [12, 13, 14]). However, they are often limited to coarse sampling (e.g., spaced at $5/10$at.%) due to the combinatorial complexity in number of alloys and low-dimensional points (e.g., $d=4$) due to the combinatorial complexity in component interactions tracked in CALPHAD calculations increasing single evaluation cost [10, 11], sometimes limited further to particular points such as equimolar alloys [15].
@@
To somewhat alleviate these computational cost challenges, ML models started to be used as surrogates for thermodynamic calculations and experiments [16, 17] or in the role of candidate selection from ML latent space [18]. They are extremely fast relative to traditional methods, usually taking microseconds per prediction, and they may seem to work near-instantly when used as a drop-in replacement. However, when one tries to deploy ML models on more complex systems, the combinatorial complexities involved (discussed in Section 1.4) may quickly make ML deployment very expensive, prompting optimization of the approach.
@@
While the ML inference is typically optimized to the computational limits in state-of-the-art tools like PyTorch [19], the rest of the customized composition space infrastructure, traditionally designed for thousands of evaluations taking seconds, may become a severe bottleneck when moving to billions of evaluations taking microseconds, as explored throughout this paper. In particular, being able to do the following tasks in nanosecond to microsecond region typically becomes critical and needs to be considered:
@@
    Efficient random sampling from the uniform grids and continuous distributions (Monte Carlo in Section 2.1) to facilitate approaches including active learning [18] and generative design [20].
@@ 1.3 Path Planning in Functionally Graded Materials
Another class of materials where complex compositional spaces have to be considered, even if intermediate compositions may not be complex themselves, is the class of Functionally Graded Materials (FGMs), sometimes narrowed to Compositionally Graded Materials (CGMs). In them, a set of compositions is traversed to form a compositional path inside a single physical part in order to spatially leverage combinations of properties that may not be possible or feasible with a homogeneous material [21]. In the simplest binary example, this could mean increasing the porosity fraction as a function of depth from the part surface to achieve a higher performance-to-weight ratio.
This paper focuses on the computational design of advanced FGMs, which enable solutions to otherwise impossible challenges. An example of such is the design of compositional pathways between stainless steel and titanium alloys to allow for additive manufacturing (AM) of aerospace and nuclear components, combining these alloys within a single print [22]. Such a task is highly non-trivial as the simple linear mixing causes several brittle or otherwise detrimental Fe-Ti and Cr-Ti phases to form, regardless of how gradual the compositional gradient is [23]. Formation of such phases in significant quantities is not specific to this alloy pair; thus, all possible ternary systems in Cr-Fe-Ni-Ti-V space had to be considered and manually arranged together by experts to obtain a pathway navigating through feasible regions [22].
@@
While in recent years, the fabrication of FGMs has become dominated by Directed Energy Deposition AM for various geometries (e.g., radial deposition [24]), several other notable manufacturing techniques allow the deployment of such pathways. These include deposition-based methods for high-precision applications, casting-based methods for high-volume fabrication [21], and recently, brazing consecutive metallic foils [25] to create relatively thin compositionally graded interfaces on mass.
In a typical FGM manufacturing scenario, a discrete set of compositions (individual available materials) exists in a compositional (simplex) space formed by a union of all components (usually chemical elements or compounds - not affecting later steps), as depicted in the top of Figure 1, which one could call the elemental space. The position in this elemental space is fundamental and is usually the one considered in both mechanistic (e.g., thermodynamic CALPHAD-type models [26]) and predictive (ML/empirical-rule) modeling. However, during the FGM design, it is more convenient to consider another compositional space formed by treating the original available compositions as components, as depicted on the bottom of Figure 1, which one could call attainable compositions space or more generally the design space.
Within an FGM manufacturing apparatus, it is common for each of the available compositions to be treated equally, e.g., powder hoppers [27], sputtering targets [25], or other flow sources are symmetrically arranged and offer the same degree of control. Thus, as depicted in Figure 1, the attainable compositional space can be treated as a standard simplex for design purposes and partitioned equally across dimensions, reflecting the nature of the problem even though equidistant points in it may not be equidistant in the original (elemental) space.
@@
A clear and industry-significant example of the above happens during FGM design in elemental spaces containing Hf and Zr. The two are very difficult to separate, causing both price and greenhouse emissions to rise sharply as a function of the separation purity requirements. Furthermore, physical form factors available from suppliers tend to be limited or at lower demand for pure-Zr and pure-Hf, furthering the cost. In the case of AM using wires as feedstock (WAAM) [28], as explored in detail in Appendix A, using pure Zr in place of the more industry-common alloy with $4.5\%$Hf can be somewhere from a few times to over 100 times more expensive. In a typical, manual FGM design, a researcher selects one of the two grades based on their expertise. However, by considering the two grades as independent components of the higher-dimensional design space, one can avoid forcing a decision before exploring the space, thus limiting human bias and allowing exploration of both options simultaneously, allowing their combination where regions of space insensitive to the Hf content utilize the cheaper grade while the pure Zr is used when necessary or favorable based on some path heuristic.
@@
With the design space carefully set up, one can start to evaluate different paths across it. Typically, the core considerations deal with meeting specific feasibility (hard) constraints. In the case of metal alloy FGMs, these can be (1) formation of detrimental phases based on thermodynamic equilibrium [27], (2) formation of detrimental phases based on non-equilibrium predictions of solidification results based on Scheil–Gulliver method, which better describes the as-made material [29], or (3) a combination of the two [22]. In the future, these will likely be extended through (4) precipitation modeling improving the metastable material design, thanks to the recent release of open-source high-performance software Kawin [30], and (5) automated modeling of manufacturing constraints, such as printability in AM [31]. Furthermore, one can also try to meet desirability (soft) constraints, such as the physical appearance of a composite material, which can be broken if needed. These two types of constraints are depicted in Figure 2, alongside an example path navigating through them.
@@
In addition to the feasibility and desirability constraints, further considerations are often made to how the path optimizes values of a set of properties of interest, either individually or through some heuristics combining them. Usually, this optimization constitutes finding the path that minimizes or maximizes either average or extreme values over the set of visited states, exemplified by the pink path in Figure 3. In the case of metal alloy FGMs, this can mean, for instance, minimizing the average evaporation rate of the molten metal [32], minimizing the maximum susceptibility to different cracking mechanisms [33], or maximizing the ductility [34].
Figure 3: Two path examples in the attainable compositional space annotated with some virtual property. One (pink/inner) minimizes/maximizes the average property value given a number of fixed path lengths, and another (purple/outer) minimizes the gradient in the property along the path.
The last, fundamentally different, property optimization task has to do with the gradient, or more generally, the character, of transitions between intermediate states, which will be critical later in the context of graphs in Section 4. Most commonly, one optimizes the path to minimize value function gradients, exemplified by the purple path in Figure 3, in order to, for instance, minimize the thermal expansion coefficient mismatch and by extension stresses induced by temperature changes [35].
@@ 1.4 Combinatorial Complexities
The second equation gives the number of ways $n_{d}$ balls can be arranged in $d$ bins, which is well known to be equivalent to much simpler problems of choosing $d-1$ moves or $n_{d}$ placements from $d-1+n_{d}$ possible options (see [36] or [37]). While these may seem unrelated to compositions, the former problem is precisely equivalent to finding a composition of an integer or distributing $n_{d}$ compositional fractions $\frac{1}{n_{d}}$ across components or chemical elements, giving the number $N_{S}$ of unique discrete compositions in the simplex space.
@@ 2.1 Monte Carlo
Software (e.g., alchemyst/ternplot in Matlab [38]) and methods dealing with low-dimensional or otherwise small compositional spaces, often utilize a naïve approach of sampling uniformly distributed points from a Cartesian space/grid in $d-1$ dimensions and then rejecting some infeasible points ($\sum^{d}_{i}x_{i}>1$), as depicted in the left part of Figure 4, which for small ($d\leq 4$) can be both easiest and computationally fastest.
However, this method becomes inefficient for large $d$ because the fraction of rejected points increases with the dimensionality. While this problem is widely noted in the literature [39], best to the authors’ knowledge, it has yet to be discussed quantitatively despite being critical to estimating the sampling’s computational complexity. Thus, it is derived herein.
@@
To circumvent the rejection problem, one may randomly sample from N-cube and normalize to 1; however, as shown in the center of Figure 4 and commonly known in the literature [40], this leads to oversampling in the center of each underlying dimension.
Thus, to achieve a uniform random sampling, nimplex and other carefully designed methods (e.g., [39] and [40]) tend to take Dirichlet distribution, where one samples points y from Gamma distributions with density $\frac{y_{i}^{\alpha-1}e^{-y_{i}}}{\Gamma(\alpha)}$ and consider its special "flat" case, where $\alpha=1$ simplifies the density equation to just $\frac{1e^{-y_{i}}}{1}=e^{-y_{i}}$. This is equivalent to sampling z from linear distributions and calculating $y_{i}=-\log(z_{i})$, which then can be normalized to obtain x as $x_{i}=y_{i}/\sum\textbf{y}$. The following snippet shows nimplex’s implementation of this, which samples z with the high-performance xoroshiro128+ random number generator [41] underlying randomTensor function from the Arraymancer tensor library [42].
proc simplex_sampling_mc(
dim: int, samples: int): Tensor[float] =
let neglograndom =
randomTensor[float]([samples, dim], 1.0
).map(x => -ln(x))
let sums = neglograndom.sum(axis=1)
return neglograndom /. sums
An alternative approach worth mentioning, sometimes found in this context, is based on (1) generating a $(d+1)$-length list composed of 00, $d-1$ random numbers, and $1$, (2) sorting it, and (3) obtaining $d$-length list of differences between consecutive elements, which is guaranteed to be uniformly distributed over a simplex as shown in [43]. While this approach may be easier to conceptualize, it is much more computationally expensive due to the sorting step. On the author’s laptop, for $d=9$, the method implemented in nimplex (involving calculation of 9 logarithms and normalizing them) takes $3.6$ns while the above (implemented with merge sort) takes $74.5$ns per iteration, i.e., over 20 times longer while not providing any clear benefit. Furthermore, their complexities are $\mathcal{O}(N)$ and $\mathcal{O}(N\ln N)$, respectively, so the computational cost difference will also slowly widen with increasing $d$.

### v2

@@
Efficient generation of grids and traversal graphs in compositional spaces towards exploration and path planning
@@ 1.1 Compositional Spaces
This work introduces a new high-performance library called nimplex or NIM library for simPLEX spaces, created exclusively for working with such spaces efficiently and qualitatively correctly. It was written in low-level Nim language, allowing for careful optimizations without sacrificing readability, and then compiled with a native Python interface for general use. nimplex provides an efficient implementation of (a) several existing methods from literature, discussed in Sections 2.1 and 3.1), (b) modifications of existing methods discussed in Section 3.2), and (c) entirely new capability of representing compositional spaces through graphs enabled by novel algorithm, progressively developed over the course of Section 4. Furthermore, Section 4.4 shows how nimplex can be leveraged to tackle compositional design problems in innovative ways.
Neither compositional spaces nor nimplex are exclusive to any discipline; however, to better showcase its capabilities, two complex, highly-dimensional materials-related problems of high impact are highlighted. At the same time, translating them and their solutions to other areas can be done directly effortlessly and is discussed throughout this work.
@@ 1.2 Compositionally Complex Materials
An exemplar of how tackling highly-dimensional problems allows researchers to unlock novel solutions is the class of Compositionally Complex Materials (CCMs), which combine several components (e.g., elements, compounds, or monomers) at significant fractions to take advantage of emergent behaviors, often enabled by the high configurational entropy. CCMs include several sub-classes, such as inorganic Multi Principle Element Alloys (MPEAs) or High Entropy Alloys (HEAs) [Yeh2004NanostructuredOutcomes, Cantor2004MicrostructuralAlloys, Senkov2019HighAlloys, Li2019MechanicalAlloys, Debnath2021GenerativeAlloys], High Entropy Ceramics (HECs) [Oses2020High-entropyCeramics], and High Entropy Metallic Glasses (HEMGs) [Wang2014High-EntropyGlasses], or organic High Entropy Polymers (HEPs) [Huang2021High-entropySeparation, Hou2021Entropy-DrivenPolymers, Hirai2022High-entropyReaction], and complex biomolecular condensates [Jacobs2021Self-AssemblyComponents]. Furthermore, CCMs are conceptually equivalent to a number of systems found in disciplines seemingly unrelated to materials, such as polypharmacy in the context of study of drug interactions [Chou2006TheoreticalStudies, Maher2014ClinicalElderly, Guthrie2015The19952010], or complex microbial communities, which can be synthesized to study microbial dynamics [Wu2024Data-drivenCommunities, vandenBerg2022EcologicalCommunities, vanLeeuwen2023SyntheticApplications], and methods introduced in this work should be directly applicable to them as well.
@@
Out of all CCMs, HEAs are the most extensively studied class, with thousands of literature publications spanning decades since two pioneering 2004 works by Yeh et al. [Yeh2004NanostructuredOutcomes] and by Cantor et al. [Cantor2004MicrostructuralAlloys], who independently proposed that equimolar (equal fractions) metallic alloys with more than 5 (Yeh) or between 6 and 9 (Cantor) elements, could form single solid solutions (SSS) thanks to the configurational entropy stabilizing them. Other notable HEA definitions include all alloys with idealized configurational entropy $\Delta S_{conf}\geq R\ln{5}=1.61R$ [Li2019MechanicalAlloys] ($\approx 2.32$ bits of information in the composition x) or $\Delta S_{conf}\geq 1R$ [Senkov2019HighAlloys] ($\approx 1.44$ bits).
@@
Regardless of the exact class or definition, while individual CCMs contain a few components, they always occupy very high dimensional problem spaces relative to other materials because they are not as restricted in terms of which elements are present. This results in homogeneous datasets often occupying over 30-dimensional spaces (or 10-20 for specific problems, like refractory HEA [Senkov2019HighAlloys]), which are orders of magnitude larger compared to traditional materials with one or two primary elements. This introduces opportunities for finding exceptional candidates in little-explored chemical spaces, as demonstrated by some HEAs with excellent hardness [Senkov2010RefractoryAlloys], ductility [Zhang2019PrecipitationAlloy], room temperature strength [Long2019AProperties], and refractory strength [Senkov2016DevelopmentSuperalloy, Kang2021SuperiorProcess].
@@
In recent years, high-throughput thermodynamics-driven combinatorial studies on CCMs have been successfully performed to generate high-performance materials[Elder2023ComputationalValidation, Elder2023ComputationalDown-selection], utilizing CALPHAD thermodynamic databases for CCMs/HEAs (e.g., [Ostrowska2020ThermodynamicW, Ostrowska2022ThermodynamicExperiments, Gambaro2024CombinedAlloys]). However, they are often limited to (a) coarse sampling (e.g., spaced at $5/10$at.or specific ratios) due to the combinatorial complexity in number of alloys, (b) low-dimensional points (e.g., $d=4$) due to the combinatorial complexity in component interactions tracked in CALPHAD calculations and increasing single evaluation cost [Elder2023ComputationalValidation, Elder2023ComputationalDown-selection], or sometimes even (c) limited further to particular points such as equimolar alloys [Yan2021AcceleratedLearning].
@@
To somewhat alleviate these computational cost challenges, ML models started to be used as surrogates for thermodynamic calculations and experiments [Debnath2023ComparingAlloys, Tandoc2023MiningAlloys] or in the role of candidate selection from ML latent space [Rao2022MachineDiscovery]. They are extremely fast relative to traditional methods, usually taking microseconds per prediction, and they may seem to work near-instantly when used as a drop-in replacement. However, when one tries to deploy ML models on more complex systems, the combinatorial complexities involved (discussed in Section 1.4) may quickly make ML deployment very expensive, prompting optimization of the approach.
@@
While the ML inference is typically optimized to the computational limits in state-of-the-art tools like PyTorch [Paszke2019PyTorch:Library], the rest of the customized composition space infrastructure, traditionally designed for thousands of evaluations taking seconds, may become a severe bottleneck when moving to billions of evaluations taking microseconds, as explored throughout this paper. In particular, being able to do the following tasks in nanosecond to microsecond region typically becomes critical and needs to be considered:
@@
    Efficient random sampling from the uniform grids and continuous distributions (Monte Carlo in Section 2.1) to facilitate approaches including active learning [Rao2022MachineDiscovery] and generative design [Debnath2021GenerativeAlloys].
@@ 1.3 Path Planning in Functionally Graded Materials
Another class of materials where complex compositional spaces have to be considered, even if intermediate compositions may not be complex themselves, is the class of Functionally Graded Materials (FGMs), sometimes narrowed to Compositionally Graded Materials (CGMs). In them, a set of compositions is traversed to form a compositional path inside a single physical part in order to spatially leverage combinations of properties that may not be possible or feasible with a homogeneous material [Saleh202030Challenges]. In the simplest binary example, this could mean increasing the porosity fraction as a function of depth from the part surface to achieve a higher performance-to-weight ratio [Chen2023FunctionallyReview].
The computational design of advanced FGMs enable solutions to otherwise impossible challenges, such as the creation of compositional pathways between stainless steel and titanium alloys to allow for additive manufacturing (AM) of aerospace and nuclear components, combining these alloys within a single print [Bobbio2022DesignCompositions]. Such a task is highly non-trivial as the simple linear mixing causes several brittle or otherwise detrimental Fe-Ti and Cr-Ti phases to form, regardless of how gradual the compositional gradient is [Reichardt2016DevelopmentManufacturing]. Formation of such phases in significant quantities is not specific to this alloy pair; thus, all possible ternary systems in Cr-Fe-Ni-Ti-V space had to be considered and manually arranged together by experts to obtain a pathway navigating through feasible regions [Bobbio2022DesignCompositions].
While in recent years, the fabrication of FGMs has become dominated by Directed Energy Deposition AM for various geometries (e.g., radial deposition [Hofmann2014DevelopingManufacturing]), several other notable manufacturing techniques allow the deployment of such pathways. These include deposition-based methods for high-precision applications, casting-based methods for high-volume fabrication [Saleh202030Challenges], and recently, brazing consecutive metallic foils [Wu2023ATechnology] to create relatively thin compositionally graded interfaces on mass.
In a typical FGM manufacturing scenario, a discrete set of compositions (individual available materials) exists in a compositional (simplex) space formed by a union of all components (usually chemical elements or compounds - not affecting later steps), as depicted in the top of Figure 1, which one could call the elemental space. The position in this elemental space is fundamental and is usually the one considered in both mechanistic (e.g., thermodynamic CALPHAD-type models [Olson2023GenomicDynamics]) and predictive (ML/empirical-rule) modeling. However, during the FGM design, it is more convenient to consider another compositional space formed by treating the original available compositions as components, as depicted on the bottom of Figure 1, which one could call attainable compositions space or more generally the design space.
@@
Within an FGM manufacturing apparatus, it is common for each of the available compositions to be treated equally, e.g., powder hoppers [Reichardt2021AdvancesMaterials], sputtering targets [Wu2023ATechnology], or other flow sources are symmetrically arranged and offer the same degree of control. Thus, as depicted in Figure 1, the attainable compositional space can be treated as a standard simplex for design purposes and partitioned equally across dimensions, reflecting the nature of the problem even though equidistant points in it may not be equidistant in the original (elemental) space. Notably, a critical consequence of such a setup is the ability to control the relative spacing of the design grid in different dimensions of the elemental space to enable careful control of the element ranges. For instance, one can consider an attainable space constructed from several individual metallic alloys and their equal mixture with 3% of carbon added, enabling fine control of its content in the design.
@@
A clear and industry-significant example of the above happens during FGM design in elemental spaces containing Hf and Zr. The two are very difficult to separate, causing both price and greenhouse emissions to rise sharply as a function of the separation purity requirements. Furthermore, physical form factors available from suppliers tend to be limited or at lower demand for pure-Zr and pure-Hf, furthering the cost. In the case of AM using wires as feedstock (WAAM) [Shen2016FabricationProcess], as explored in detail in Appendix A, using pure Zr in place of the more industry-common alloy with $4.5\%$Hf can be somewhere from a few times to over 100 times more expensive. In a typical, manual FGM design, a researcher selects one of the two grades based on their expertise. However, by considering the two grades as independent components of the higher-dimensional design space, one can avoid forcing a decision before exploring the space, thus limiting human bias and allowing exploration of both options simultaneously, allowing their combination where regions of space insensitive to the Hf content utilize the cheaper grade while the pure Zr is used when necessary or favorable based on some path heuristic.
With the design space carefully set up, one can start to evaluate different paths across it. Typically, the core considerations deal with meeting specific feasibility (hard) constraints. In the case of metal alloy FGMs, these can be (1) formation of detrimental phases based on thermodynamic equilibrium [Reichardt2021AdvancesMaterials], (2) formation of detrimental phases based on non-equilibrium predictions of solidification results based on Scheil–Gulliver method, which better describes the as-made material [Bocklund2020ExperimentalMaterials], or (3) a combination of the two [Bobbio2022DesignCompositions]. In the future, these will likely be extended through (4) precipitation modeling improving the metastable material design, thanks to the recent release of open-source high-performance software Kawin [Ury2023Kawin:Model], and (5) automated modeling of manufacturing constraints, such as printability in AM [SheikhAnAlloys]. Furthermore, one can also try to meet desirability (soft) constraints, such as the physical appearance of a composite material, which can be broken if needed. These two types of constraints are depicted in Figure 2, alongside an example path navigating through them.
@@
In addition to the feasibility and desirability constraints, further considerations are often made to how the path optimizes values of a set of properties of interest, either individually or through some heuristics combining them. Usually, this optimization constitutes finding the path that minimizes or maximizes either average or extreme values over the set of visited states, exemplified by the pink path in Figure 3. In the case of metal alloy FGMs, this can mean, for instance, minimizing the average evaporation rate of the molten metal [Mukherjee2016PrintabilityManufacturing], minimizing the maximum susceptibility to different cracking mechanisms [Yang2023DesignCracking], or maximizing the ductility [Hu2021ScreeningAlloys].
Figure 3: Two path examples in the attainable compositional space overlaid with hypothetical property values, for illustration purposes. One (pink/inner) minimizes/maximizes the average property value given a number of fixed path lengths, and another (purple/outer) minimizes the gradient in the property along the path.
The last, fundamentally different, property optimization task has to do with the gradient, or more generally, the character, of transitions between intermediate states, which will be critical later in the context of graphs in Section 4. Most commonly, one optimizes the path to minimize value function gradients, exemplified by the purple path in Figure 3, in order to, for instance, minimize the thermal expansion coefficient mismatch and by extension stresses induced by temperature changes [Kirk2021ComputationalMonotonicity].
@@ 1.4 Combinatorial Complexities
The second equation gives the number of ways $n_{d}$ balls can be arranged in $d$ bins, which is well known to be equivalent to much simpler problems of choosing $d-1$ moves or $n_{d}$ placements from $d-1+n_{d}$ possible options (see [Nijenhuis1978CombinatorialCalculators] or [Chasalow1995AlgorithmPoints]). While these may seem unrelated to compositions, the former problem is precisely equivalent to finding a composition of an integer or distributing $n_{d}$ compositional fractions $\frac{1}{n_{d}}$ across components or chemical elements, giving the number $N_{S}$ of unique discrete compositions in the simplex space.
@@ 2.1 Monte Carlo
Software (e.g., alchemyst/ternplot in Matlab [Sandrock2017Alchemyst/ternplothttps://github.com/alchemyst/ternplot]) and methods dealing with low-dimensional or otherwise small compositional spaces, often utilize a naïve approach of sampling uniformly distributed points from a Cartesian space/grid in $d-1$ dimensions and then rejecting some infeasible points ($\sum^{d}_{i}x_{i}>1$), as depicted in the left part of Figure 4, which for small ($d\leq 4$) can be both easiest and computationally fastest.
However, this method becomes inefficient for large $d$ because the fraction of rejected points increases with the dimensionality. While this problem is widely noted in the literature [Allen2022AAlloys], best to the authors’ knowledge, it has yet to be discussed quantitatively despite being critical to estimating the sampling’s computational complexity. Thus, it is derived herein.
@@
To circumvent the rejection problem, one may randomly sample from N-cube and normalize to 1; however, as shown in the center of Figure 4 and commonly known in the literature [Otis2017AnSystems], this leads to oversampling in the center of each underlying dimension.
Thus, to achieve a uniform random sampling, nimplex and other carefully designed methods (e.g., [Allen2022AAlloys] and [Otis2017AnSystems]) tend to take Dirichlet distribution, where one samples points y from Gamma distributions with density $\frac{y_{i}^{\alpha-1}e^{-y_{i}}}{\Gamma(\alpha)}$ and consider its special "flat" case, where $\alpha=1$ simplifies the density equation to just $\frac{1e^{-y_{i}}}{1}=e^{-y_{i}}$. This is equivalent to sampling z from linear distributions and calculating $y_{i}=-\log(z_{i})$, which then can be normalized to obtain x as $x_{i}=y_{i}/\sum\textbf{y}$. The following snippet shows nimplex’s implementation of this, which samples z with the high-performance xoroshiro128+ random number generator [Blackman2018ScrambledGenerators] underlying randomTensor function from the Arraymancer tensor library [RatsimbazafyMratsim/Arraymancer:Backends].
{minted}
[linenos=true, mathescape]nim proc simplex_sampling_mc( dim: int, samples: int): Tensor[float] = let neglograndom = randomTensor[float]([samples, dim], 1.0 ).map(x => -ln(x)) let sums = neglograndom.sum(axis=1) return neglograndom /. sums
An alternative approach worth mentioning, sometimes found in this context, is based on (1) generating a $(d+1)$-length list composed of 00, $d-1$ random numbers, and $1$, (2) sorting it, and (3) obtaining $d$-length list of differences between consecutive elements, which is guaranteed to be uniformly distributed over a simplex as shown in [Rubin1981TheBootstrap]. While this approach may be easier to conceptualize, it is much more computationally expensive due to the sorting step. On the author’s laptop, for $d=9$, the method implemented in nimplex (involving calculation of 9 logarithms and normalizing them) takes $3.6$ns while the above (implemented with merge sort) takes $74.5$ns per iteration, i.e., over 20 times longer while not providing any clear benefit. Furthermore, their complexities are $\mathcal{O}(N)$ and $\mathcal{O}(N\ln N)$, respectively, so the computational cost difference will also slowly widen with increasing $d$.